OpenStudy (anonymous):

7/8-1/x=3/4 find x would x be 4?

OpenStudy (anonymous):

\[\frac{ 7 }{ 8 } - \frac{ 1 }{ x } = \frac{ 3 }{ 4 }\]

OpenStudy (anonymous):

\[\frac{ 7 }{ 8}-\frac{ 1 }{ x}=\frac{ 3}{ 4 }\] First, let's make the denominators (other than x) the same; \[4\times2=8\] So what you do to the denominator you do to the numerator; \[3\times2=6\] Which makes it; \[\frac{ 7 }{ 8 }-\frac{ 1 }{ x }=\frac{ 6 }{ 8 }\] So, for this to be true, 7-1=6, but for that to happen the denominators have to be the same, which would mean; \[x=8\]